How would I perform this indicated operation?
5x^2y^3(7x^3 – 8x^2y^2 + 9xy^4)

Answer by CharlesG2(834)

How would I perform this indicated operation?
5x^2y^3(7x^3 – 8x^2y^2 + 9xy^4)
by distributive property --> a(b + c + d) = ab + ac + ad
and you add exponents when multiplying --> x^2 * x^3 = x^(2+3) = x^5
5x^2y^3 * 7x^3 - 5x^2y^3 * 8x^2y^2 + 5x^2y^3 * 9xy^4
35x^5y^3 - 40x^4y^5 + 45x^3y^7